# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

### v3

@@ V Conclusion
A novel sum-rate maximization scheme for RBD-RIS-aided MU-MISO systems is presented. To this end, an optimization problem was formulated to determine the BD-RIS scattering matrix configuration that maximizes the users’ sum-rate, and solved using a Riemannian manifold steepest descent method with a closed-form gradient. Simulation results validated the effectiveness of the proposed approach, showing consistent performance gains over the SotA, where the improvements arise from jointly exploiting the geometric structure of the manifold and the direct sum-rate based problem formulation. Overall, the results show that manifold-based scattering matrix optimization enables scalable and robust performance in RIS-aided systems. Future work will extend the framework to include transmit beamforming design.
@@ Appendix A Proof of Lemma 1
$\mathbf{A}=\begin{pmatrix}a_{1,1}&a_{1,2}&\cdots&a_{1,N}\\ a_{2,1}&a_{2,2}&\cdots&a_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N,1}&a_{N,2}&\cdots&a_{N,N}\end{pmatrix}$. (46)
@@
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},$ (47)
@@
Proof.
For simplicity, let $\phi=\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}$. Then,
$\frac{\partial\phi}{\partial\Re{\mathbf{A}}}$ $=\begin{pmatrix}\frac{\partial\phi}{\partial\Re a_{1,1}}&\frac{\partial\phi}{\partial\Re a_{1,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{1,N}}\\ \frac{\partial\phi}{\partial\Re a_{2,1}}&\frac{\partial\phi}{\partial\Re a_{2,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Re a_{N,1}}&\frac{\partial\phi}{\partial\Re a_{N,2}}&\cdots&\frac{\partial\phi}{\partial\Re a_{N,N}}\end{pmatrix}$ (49)
$=2\Re{\begin{pmatrix}a_{1,1}&a_{2,1}&\cdots&a_{N,1}\\ a_{1,2}&a_{2,2}&\cdots&a_{N,2}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,N}&a_{2,N}&\cdots&a_{N,N}\end{pmatrix}}$
@@
$j\frac{\partial\phi}{\partial\Im{\mathbf{A}}}$ $=\begin{pmatrix}\frac{\partial\phi}{\partial\Im a_{1,1}}&\frac{\partial\phi}{\partial\Im a_{1,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{1,N}}\\ \frac{\partial\phi}{\partial\Im a_{2,1}}&\frac{\partial\phi}{\partial\Im a_{2,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Im a_{N,1}}&\frac{\partial\phi}{\partial\Im a_{N,2}}&\cdots&\frac{\partial\phi}{\partial\Im a_{N,N}}\end{pmatrix}$ (50)
$=2j\Im{\begin{pmatrix}a_{1,1}&a_{2,1}&\cdots&a_{N,1}\\ a_{1,2}&a_{2,2}&\cdots&a_{N,2}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,N}&a_{2,N}&\cdots&a_{N,N}\end{pmatrix}}$,
where it is clear that
$\frac{\partial\phi}{\partial\Re\mathbf{A}}=2\Re\mathbf{A}^{T},$ (51)
and
$j\frac{\partial\phi}{\partial\Im\mathbf{A}}=2j\Im\mathbf{A}^{T}.$ (52)
Straightforwardly, the Generalized Complex Derivative and Conjugate Complex Derivative [38] are obtained as
@@
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i}^{N}|a_{i,i}|^{2}+2\Re\{\sum_{i}^{N}\sum_{k=i+1}^{N}a_{i,k}^{\star}a_{k,i}\},$ (55)
as such, the Complex Gradient Matrix [38] can be computed as follows
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},$ (56)
hence, (24) is proven.
@@ Appendix B Proof of Lemma 2
$d=\left(\!\begin{array}[]{c}a_{1}\\ a_{2}\\ \vdots\\ a_{N}\end{array}\!\right)^{T}\!\!\!\left(\!\begin{array}[]{cccc}B_{1,1}&B_{1,2}&\cdots&B_{1,N}\\ B_{2,1}&B_{2,2}&\cdots&B_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ B_{N,1}&B_{N,2}&\cdots&B_{N,N}\end{array}\!\right)\!\!\!\left(\!\begin{array}[]{c}c_{1}\\ c_{2}\\ \vdots\\ c_{N}\end{array}\!\right).$ (57)
@@
$\nabla_{\mathbf{B}}|d|^{2}=d\mathbf{a}^{\star}\mathbf{c}^{H},$ (58)
@@
Following the steps for calculating the gradient of complex valued functions, as described in [38]. The derivatives with respect to the real and imaginary parts of $\mathbf{B}$ are calculated, as shown below in
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Re\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}},$ (59)
@@
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Im\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}.$ (60)
@@
$\frac{\partial|d|^{2}}{\partial d}=d^{\star},$ (61)
$\frac{\partial|d|^{2}}{\partial d^{\star}}=d.$ (62)
@@
$\frac{\partial d}{\partial\Re\mathbf{B}}$ $=\begin{pmatrix}\frac{\partial d}{\partial\Re B_{1,1}}&\frac{\partial d}{\partial\Re B_{1,2}}&\cdots&\frac{\partial d}{\partial\Re B_{1,N}}\\ \frac{\partial d}{\partial\Re B_{2,1}}&\frac{\partial d}{\partial\Re B_{2,2}}&\cdots&\frac{\partial d}{\partial\Re B_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Re B_{N,1}}&\frac{\partial d}{\partial\Re B_{N,2}}&\cdots&\frac{\partial d}{\partial\Re B_{N,N}}\end{pmatrix}$ (63)
$={\begin{pmatrix}a_{1}c_{1}&a_{1}c_{2}&\cdots&a_{1}c_{N}\\ a_{2}c_{1}&a_{2}c_{2}&\cdots&a_{2}c_{N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N}c_{1}&a_{N}c_{2}&\cdots&a_{N}c_{N}\end{pmatrix}}=\mathbf{a}\mathbf{c}^{T},$
$\frac{\partial d}{\partial\Im\mathbf{B}}$ $=\begin{pmatrix}\frac{\partial d}{\partial\Im B_{1,1}}&\frac{\partial d}{\partial\Im B_{1,2}}&\cdots&\frac{\partial d}{\partial\Im B_{1,N}}\\ \frac{\partial d}{\partial\Im B_{2,1}}&\frac{\partial d}{\partial\Im B_{2,2}}&\cdots&\frac{\partial d}{\partial\Im B_{2,N}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Im B_{N,1}}&\frac{\partial d}{\partial\Im B_{N,2}}&\cdots&\frac{\partial d}{\partial\Im B_{N,N}}\end{pmatrix}$ (64)
$=j{\begin{pmatrix}a_{1}c_{1}&a_{1}c_{2}&\cdots&a_{1}c_{N}\\ a_{2}c_{1}&a_{2}c_{2}&\cdots&a_{2}c_{N}\\ \vdots&\vdots&\ddots&\vdots\\ a_{N}c_{1}&a_{N}c_{2}&\cdots&a_{N}c_{N}\end{pmatrix}}=j\mathbf{a}\mathbf{c}^{T}.$
@@
$\frac{\partial d^{\star}}{\partial\Re\mathbf{B}}=\mathbf{a}^{\star}\mathbf{c}^{H},$ (65)
$\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}=-j\mathbf{a}^{\star}\mathbf{c}^{H}.$ (66)
Finally
@@
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=jd^{\star}\mathbf{a}\mathbf{c}^{T}-jd\mathbf{a}^{\star}\mathbf{c}^{H}.$ (68)
$\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=\frac{1}{2}\Bigg(\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}+j\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}\Bigg)=d\mathbf{a}^{\star}\mathbf{c}^{H}.$ (69)
@@
$\nabla_{\mathbf{B}}|d|^{2}=2\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=2d\mathbf{a}^{\star}\mathbf{c}^{H}=2\big(d^{\star}\mathbf{a}\mathbf{c}^{T}\big)^{\star},$ (70)

### v4

@@ V Conclusion
A novel sum-rate maximization scheme for RBD-RIS-aided MU-MISO systems is presented. To this end, an optimization problem was formulated to determine the BD-RIS scattering matrix configuration that maximizes the users’ sum-rate, and solved using a Riemannian manifold steepest descent method with a closed-form gradient. Simulation results validated the effectiveness of the proposed approach, showing consistent performance gains over the SotA, where the improvements arise from jointly exploiting the geometric structure of the manifold and the direct sum-rate based problem formulation. Additionally, the simulation results were compared against joint scattering matrix and beamforming designs. Under the considered scenarios, the proposed method outperformed the design with fixed beamforming, which opens up the possibility for beamforming optimization.
This is promising, as further beamforming design optimization is expected to provide even greater performance improvements. Overall, the results show that manifold-based scattering matrix optimization enables scalable and robust performance in RIS-aided systems. Future work will extend the framework to include transmit beamforming design.
@@ Appendix A Proof of Lemma 1
[…]
@@
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},\vskip-4.30554pt$ (48)
@@
Proof: For simplicity, let $\phi=\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}$. Then,
$\frac{\partial\phi}{\partial\Re{\mathbf{A}}}={\begin{pmatrix}\frac{\partial\phi}{\partial\Re a_{1,1}}&\cdots&\frac{\partial\phi}{\partial\Re a_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Re a_{N,1}}&\cdots&\frac{\partial\phi}{\partial\Re a_{N,N}}\end{pmatrix}=2\Re\mathbf{A}^{T},}\vskip-4.30554pt$ (50)
@@
$j\frac{\partial\phi}{\partial\Im{\mathbf{A}}}={\begin{pmatrix}\frac{\partial\phi}{\partial\Im a_{1,1}}&\cdots&\frac{\partial\phi}{\partial\Im a_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial\phi}{\partial\Im a_{N,1}}&\cdots&\frac{\partial\phi}{\partial\Im a_{N,N}}\end{pmatrix}=2j\Im\mathbf{A}^{T}.}\vskip-4.30554pt$ (51)
Straightforwardly, the Generalized Complex Derivative and Conjugate Complex Derivative [48, eq.(229), eq.(230)] are obtained as
@@
$\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=\sum_{i}^{N}|a_{i,i}|^{2}+2\Re\Big\{\sum_{i}^{N}\sum_{k=i+1}^{N}a_{i,k}^{\star}a_{k,i}\Big\},\vskip-4.30554pt$ (54)
as such, the Complex Gradient Matrix [48, eq.(233)] can be computed as follows
$\nabla_{\mathbf{A}}\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}=2\frac{\partial\mathrm{Tr}\{\mathbf{A}^{\star}\mathbf{A}\}}{\partial\mathbf{A}^{\star}}=2\mathbf{A}^{T},\vskip-4.30554pt$ (55)
hence, (30) is proven. $\square$
@@ Appendix B Proof of Lemma 2
$d={\left(\!\begin{array}[]{c}a_{1}\\[-4.30554pt] \vdots\\[-4.30554pt] a_{N}\end{array}\!\right)^{T}\!\!\!\left(\!\begin{array}[]{cccc}B_{1,1}&\cdots&B_{1,N}\\[-4.30554pt] \vdots&\ddots&\vdots\\ B_{N,1}&\cdots&B_{N,N}\end{array}\!\right)\!\!\!\left(\!\begin{array}[]{c}c_{1}\\[-4.30554pt] \vdots\\[-4.30554pt] c_{N}\end{array}\!\right).}\vskip-4.30554pt$ (56)
@@
$\nabla_{\mathbf{B}}|d|^{2}=d\mathbf{a}^{\star}\mathbf{c}^{H},\vskip-4.30554pt$ (57)
@@
Following the steps for calculating the gradient of complex valued functions, as described in [48, eq.(229), eq.(230), eq.(233)]. The derivatives with respect to the real and imaginary parts of $\mathbf{B}$ are calculated, as shown below
$\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Re\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}},\vskip-4.30554pt$ (58)
@@
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=\frac{\partial|d|^{2}}{\partial d}\frac{\partial d}{\partial\Im\mathbf{B}}+\frac{\partial|d|^{2}}{\partial d^{\star}}\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}.\vskip-4.30554pt$ (59)
@@
$\frac{\partial|d|^{2}}{\partial d}=d^{\star}\quad\text{and}\quad\frac{\partial|d|^{2}}{\partial d^{\star}}=d.$ (60)
@@
$\frac{\partial d}{\partial\Re\mathbf{B}}={\begin{pmatrix}\frac{\partial d}{\partial\Re B_{1,1}}&\cdots&\frac{\partial d}{\partial\Re B_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Re B_{N,1}}&\cdots&\frac{\partial d}{\partial\Re B_{N,N}}\end{pmatrix}=\mathbf{a}\mathbf{c}^{T},}$ (61)
$\frac{\partial d}{\partial\Im\mathbf{B}}={\begin{pmatrix}\frac{\partial d}{\partial\Im B_{1,1}}&\cdots&\frac{\partial d}{\partial\Im B_{1,N}}\\[-4.30554pt] \vdots&\ddots&\vdots\\ \frac{\partial d}{\partial\Im B_{N,1}}&\cdots&\frac{\partial d}{\partial\Im B_{N,N}}\end{pmatrix}=j\mathbf{a}\mathbf{c}^{T}.}$ (62)
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\partial d^{\star}}{\partial\Re\mathbf{B}}=\mathbf{a}^{\star}\mathbf{c}^{H}\quad\text{and}\quad\frac{\partial d^{\star}}{\partial\Im\mathbf{B}}=-j\mathbf{a}^{\star}\mathbf{c}^{H}.$ (63)
Finally we obtain
@@
$\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}=jd^{\star}\mathbf{a}\mathbf{c}^{T}-jd\mathbf{a}^{\star}\mathbf{c}^{H},$ (65)
$\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=\frac{1}{2}\Big(\frac{\partial|d|^{2}}{\partial\Re\mathbf{B}}+j\frac{\partial|d|^{2}}{\partial\Im\mathbf{B}}\Big)=d\mathbf{a}^{\star}\mathbf{c}^{H}.$ (66)
@@
$\nabla_{\mathbf{B}}|d|^{2}=2\frac{\mathrm{d}|d|^{2}}{\mathrm{d}\mathbf{B}^{\star}}=2d\mathbf{a}^{\star}\mathbf{c}^{H}=2\big(d^{\star}\mathbf{a}\mathbf{c}^{T}\big)^{\star},\vskip-4.30554pt$ (67)